If 150cm3 of steam at 100∘C is added to 250cm3 water at 40∘C, how much steam(in cm3) is condensed into water at 100∘C?
1
Expert's answer
2020-12-06T17:26:32-0500
"m_sR=m_wc(T-T_0)\\\rho_sV_sR=\rho_wV_wc(T-T_0)
\\(0.59)(2260000)V_s=(1000)(250)(4200)(100-40)\\V_s=47250\ cm^3>V_{s0}"

Thus, all steam is condensed into water:


"V_s=V_{s0}=150\ cm^3"